3. General Electric employs a job order cost accounting system and keeps perpetual inventory records. The following transactions occurred in the first month of operations: (20 pts)

1.   Direct materials requisitioned during the month:

                  Job 101                $20,000

                  Job 102                  16,000

                  Job 103                24,000

                                               $60,000

2.   Direct labor incurred and charged to jobs during the month was:

                  Job 101                $32,000

                  Job 102                  28,000

                  Job 103                20,000

                                               $80,000

3.   Manufacturing overhead was applied to jobs worked on using a predetermined overhead rate based on 75% of direct labor costs.

4.   Actual manufacturing overhead costs incurred during the month amounted to $66,000.

5.   Job 101 consisting of 1,000 units and Job 103 consisting of 200 units were completed during the month.

Instructions

Answer the following questions:

  1. How much manufacturing overhead was applied to Job 103 during the month?
  1. Compute the unit cost of Jobs 101 and 103.
  1. What is the balance in Work In Process Inventory at the end of the month?
  1. Determine if manufacturing overhead was under- or overapplied during the month. How much?
Answer #1

a) manufacturing overhead applied to Job 103 = 20000*75% = $15000

b) Calculate unit cost

Job 101 Job 103
Direct material 20000 24000
Direct labor 32000 20000
Overhead 24000 15000
Total cost 76000 59000
Unit 1000 200
Unit Cost 76 295

c) Work in process

Job 102
Direct material 16000
Direct labor 28000
Overhead 21000
Total cost 65000

d) Applied overhead = 80000*75% = 60000

actual overhead = 66000

Under applied overhead = 6000